Solve for the unknown number of years in each of the following (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.):

Present Value Years Interest Rate Future Value
$570 8 % $1,293
820 12 2,173
18,500 18 343,880
21,600 14 363,176
Answer #1

Hi,

we will use compounding formula to find number of years

first part Present Value P= $570

interest rate r = 8%

Future Value F = $1,293

Compounding Formula

F = P*(1+r)^t

1293 = 570*(1.08)^t

1.08^t = 2.27

t = log(1.08)(2.27)

t= 10.64 years

Same way for part 2

t = log(1.08)(2173/820)

t= 12.66 years

part 3

t = log(1.08)(343880/18500)

t = 37.97 years

part 4

t = log(1.08)(363176/21600)

t = 36.67 years

Thanks